OpenStudy (anonymous):

How would you find the speed of a bowling ball (assuming that the speed of sound is 340 m/s) if the ball travels with constant speed, on a 16.5 m long bowling lane, and hits the pins at 2.50 seconds after it is released from the bowler's hands?

OpenStudy (anonymous):

v=d/t

OpenStudy (ash2326):

Ball is travelling with a constant speed, distance covered 16.5 m time taken= 2.5 seconds speed= 16.5/2.5= 6.8 m/s

OpenStudy (anonymous):

I figured i would be using that formula.